SUPPOSE THAT THE PERIMETER OF A SQUARE EQUALS THE PERIMETER OF A RECTANGLE. THE WIDTH OF THE RECTANGLE IS 9 INCHES LESS THAN TWICE THE SIDE OF THE SQUARE AND THE LENGTH OF THE RECTANGLE IS 3 INCHES LESS THAN TWICE THE SIDE OF THE SQUARE. FIND THE DIMINSIONS OF THE SQUARE AND THE RECTANGLE.
Answer by venugopalramana(3286)

SUPPOSE THAT THE PERIMETER OF A SQUARE EQUALS THE PERIMETER OF A RECTANGLE. THE WIDTH OF THE RECTANGLE IS 9 INCHES LESS THAN TWICE THE SIDE OF THE SQUARE AND THE LENGTH OF THE RECTANGLE IS 3 INCHES LESS THAN TWICE THE SIDE OF THE SQUARE. FIND THE DIMINSIONS OF THE SQUARE AND THE RECTANGLE.
LET LENGTH OF SIDE =A
LET LENGTH AND WIDTH OF RECTANGLE BE L AND W
PERIMETEROF SQUARE=4A=PERIMETER OF RECTANGLE =2(L+B)
2A=L+B...................I
TWICE SIDE OF SQUARE=2A
9" LESS =2A-9=W...............II
3" LESS =2A-3=L.................III
SUBSTITUTING II AND III IN I
2A=2A-3+2A-9
2A=12
A=6
L=2A-3=2*6-3=9
W=2*6-9=3